if measure of angle A=60 degrees and side AB=5, then side BC=? It is a right triangle with the 90 degree angle at the top point of the triangle. So the triangle has A on right point of triangle, B on top point of triangle(where 90 degree angle is) and C on left point of triangle. I tried to solve by using tan, but the answer is not correct. The answer sheet says the answer is 5 sq rt 3. We haven't been taught how to do sine or cosine yet so I don't believe that will help me.
Answer by mananth(16946)

If B = 90
& A= 60
Then c=30 (A+C)=90
AB = 5
AB/BC= tan 30
AB/tan 30 = BC
5/(1/sqrt(3))=BC
BC = 5 sqrt(3)